If the dimensions of the length are expressed as Gxcyhz; where G, c and h are the universal gravitational constant, speed of light and Planck’s constant respectively, then


(a)    X = ½, y = ½


(b)    x = 1/2 , z = 1/2


(c)    y = ½, z = 3/2


        (d)    y = –3/2, z = ½


Grade:upto college level

2 Answers

Simran Bhatia
10 years ago

(b, d)

 

        L = (M–1L3T–2)x(LT–1)y(ML2T–1)z

 

        – x + z = 0, 3x + y + 2z = 1, – 2x – y – z = 0

 

Þ x = ½, y = –3/2, z = ½

Anjali kumari
5 years ago
The answer will be x=1/2 and y=-3/2 and z=1/2.Hence the answer is (B) and(D).By getting the three equations.